ALGEBRAS OVER EQUIVARIANT SPHERE SPECTRA

A. D. ELMENDORF AND J. P. MAY

Abstract. We study algebras over the sphere spectrum S_{G} of a compact Lie group G. In particular, we show how to construct S_{G}-algebras from S-algebras, where S is the nonequivariant sphere spectrum. This gives a reservoir of equivariant examples to which recently developed algebraic techniques in stable homotopy theory can be applied. A special case will be used in a companion paper of Benson and Greenlees to study the ordinary cohomology of the classifying space BG.

In [3], Kriz, Mandell, and the authors developed a theory of highly structured ring, module, and algebra spectra. Although that paper was written non-equivariantly, its first stated result was the blanket assertion that all of the rest of its general theoretical results apply verbatim to G-spectra for a compact Lie group G. We here begin to explore such highly structured equivariant spectra.

The ground category for the theory of [3] (see also [2]) is the category ℳ_{S} of S-modules, where S is the sphere spectrum; its derived category 𝒟_{S} is obtained by inverting the weak equivalences and is equivalent to the classical stable homotopy category. The ground category we are interested in is the category of S_{G}-modules, where S_{G} is the sphere G-spectrum; its derived category is equivalent to the Lewis-May equivariant stable homotopy category [7] (see also [5]).

One of the main features of the equivariant world is change of universe. A Guniverse is a countably infinite dimensional real inner product space U with an action of G through linear isometric isomorphisms such that U is the colimit of its finite dimensional representations, the trivial representation occurs in U, and each irreducible representation that occurs in U occurs infinitely often. For a universe U, a G-spectrum E indexed on U is a collection of based G-spaces EV and suitably compatible homeomorphisms EV ⟶ Ω^{W −V} EW for V ⊂ W, where W − V is the orthogonal complement of V in W.

We obtain different stable homotopy categories of G-spectra depending on which representations occur in U. If only the trivial representation occurs, then we say that U is a trivial universe and we obtain what are called naive G-spectra. In particular, any nonequivariant spectrum may be regarded as a naive G-spectrum with trivial G-action. If all irreducible representations occur, then we say that U is a complete universe and we obtain what are called genuine G-spectra, or simply G-spectra. We let G𝒮 U denote the category of G-spectra indexed on U.

We obtain a trivial universe U^{G} from a universe U by passing to fixed points. There is a change of universe functor i_{∗} : G𝒮 U^{G} ⟶ G𝒮 U associated with the inclusion i : U^{G} ⟶ U. Such change of universe functors play a central role in equivariant stable homotopy theory. We refer the reader to [7, II§1], [5, §3], or [4, I§0] for background and discussion. Recently, Lewis [6] has made a thorough study of the homotopical properties of change of universe. For example, he has proven that the stable homotopy categories of spectra indexed on universes U and U′ are equivalent if and only if the orbits G/H that embed in U are the same as the orbits that embed in U′.

Our concern is different: we are interested in obtaining point-set level change of universe functors that are so well-behaved that they preserve the highly structured ring, module, and algebra spectra introduced in [3]. This has considerable constructive power. A major gap in equivariant stable homotopy theory is that equivariant infinite loop space theory has not yet been developed for compact Lie groups: we do not have a recognition principle that allows us to construct G-spectra, let alone highly structured G-spectra, from space level data. The present theory partially rectifies this since it shows how to construct highly structured G-spectra from highly structured spectra. It is just such an example that will be needed in the companion paper of Benson and Greenlees [1].

Beyond its constructive force, our work will illuminate the general theory and drive home the importance of the obvious distinction between categories of modules and their derived categories. To avoid excessive use of subscripts, we shall let SU denote the sphere G-spectrum indexed on a G-universe U. When U is a fixed given complete G-universe, we shall write S_{G} = SU and S = SU^{G}. Thus S_{G} is our sphere G-spectrum and S is our nonequivariant sphere spectrum. However, we shall work equivariantly throughout, and we therefore regard S as a G-spectrum with trivial action by G. Let Gℳ_{SU} be the category of SU-modules and let hGℳ_{SU} be its homotopy category. Let G𝒟_{SU} be the derived category of SU-modules; it is obtained from hGℳ_{SU} by adjoining inverses to the weak equivalences, which are the maps of SU-modules that are weak equivalences of underlying G-spectra. We view the category ℳ_{SU^{G}} of nonequivariant S-modules as the full subcategory of G-trivial modules in the category Gℳ_{SU^{G}} of equivariant SU^{G}-modules.

The category G𝒟_{SU} is equivalent to the stable homotopy category of G-spectra indexed on U, and Lewis’s result shows how this category depends on U. In contrast, we shall prove the startling fact that the point-set level categories Gℳ_{SU}, and also their homotopy categories hGℳ_{SU}, are independent of U, up to natural equivalence of categories. There is no contradiction: the relevant change of universe functors do not preserve weak equivalences and so do not pass to equivalences of derived categories. The point-set level equivalences of categories preserve rings, modules, and algebras.

Recall that a functor between symmetric monoidal categories is said to be monoidal if it preserves the given products and unit objects up to isomorphisms that are suitably compatible with the respective unity, associativity, and commutativity isomorphisms. Such a functor necessarily preserves monoids, commutative monoids, and objects with actions by monoids. In the category Gℳ_{SU} of SU-modules, the monoids are the SU-algebras R and the objects with action by R are the R-modules M. These notions are defined in terms of maps of SU-modules SU ⟶ R, R ∧_{SU} R ⟶ R, and R ∧_{SU} M ⟶ M such that the usual diagrams commute. See [3, II§§3-4] for discussion and for comparison with the earlier definitions of A_{∞} and E_{∞} ring G-spectra; the discussion is given nonequivariantly, but it applies verbatim equivariantly. As a special case of more general observations, we shall prove the following formal result.

Theorem 0.1. Let U be a G-universe. There is a monoidal equivalence of categories

I_{U^{G}}^{U} : Gℳ_{SU^{G}} ⟶ Gℳ_{SU}.

Thus, if R is an SU^{G}-algebra and M is an R-module, then I_{U^{G}}^{U}R is an SU-algebra and I_{U^{G}}^{U}M is an I_{U^{G}}^{U}R-module.

When we restrict attention to homotopically well-behaved modules and suitably related universes, we find that our point-set level change of universe equivalences pass to derived categories to give new models for the homotopical change of universe functors between derived categories, such as i_{∗} : G𝒟_{SU^{G}} ⟶ G𝒟_{SU}. The details depend on the Quillen model category structures that the theory of [3] assigns to all categories in sight. We use the term “q-cofibrant” for Quillen cofibrant objects in a given category. As a special case of more general observations, we shall prove the following homotopical result.

Theorem 0.2. For all M in a certain class ℰ̄_{SU^{G}} of SU^{G}-modules, there is a natural weak equivalence of G-spectra

i_{∗}M ⟶ I_{U^{G}}^{U}M.

The class ℰ̄_{SU^{G}} contains all q-cofibrant SU^{G}-algebras, all q-cofibrant commutative SU^{G}-algebras, and all cell modules over such SU^{G}-algebras.

Combining the previous results, we conclude that I_{U^{G}}^{U} gives a point-set level model for i_{∗} that carries highly structured spectra to highly structured G-spectra. We record the example relevant to the companion paper [1], taking U to be a fixed complete G-universe and deleting it from the notations. For an S_{G}-algebra R_{G} and a G-space X, the function G-spectrum F (X_{+}, R_{G}) is an S_{G}-algebra, commutative if R_{G} is so, with product induced by the product of R_{G} and the diagonal on X [3, VII.2.10]. Let Σ_{G}^{∞} denote the suspension G-spectrum functor.

Corollary 0.3. If R is a commutative S-algebra, then F (EG_{+}, I_{U^{G}}^{U}R) is a commutative S_{G}-algebra. If R is q-cofibrant, then F (EG_{+}, I_{U^{G}}^{U}R) represents R-Borel cohomology on based G-spaces X. Precisely,

[Σ^{∞}(EG ×_{G} X)_{+}, R]^{∗} ≅ [Σ_{G}^{∞}X_{+}, F (EG_{+}, I_{U^{G}}^{U}R)]_{S_{G}}^{∗}.

Proof. On the left, we understand maps in the classical stable homotopy category. On the right, we understand maps in the derived category of S_{G}-modules, which is equivalent to the Lewis-May stable homotopy category of G-spectra; we denote maps in the latter category by [−, −]_{G}^{∗}. The first of the following isomorphisms is standard (e.g. [4, 0.3, 0.7]). The second follows from the natural isomorphism of functors i_{∗} ∘ Σ^{∞} ≅ Σ_{G}^{∞}, the commutation of the suspension G-spectrum functor with smash products with G-spaces, and adjunction. The third is a direct consequence of Theorem 0.2 and the cited equivalence of categories.

[Σ^{∞}(EG ×_{G} X)_{+}, R]^{∗} ≅ [i_{∗}Σ^{∞}(EG × X)_{+}, i_{∗}R]_{G}^{∗}
≅ [Σ_{G}^{∞}X_{+}, F (EG_{+}, i_{∗}R)]_{G}^{∗}
≅ [Σ_{G}^{∞}X_{+}, F (EG_{+}, I_{U^{G}}^{U}R)]_{S_{G}}^{∗}.

Benson and Greenlees [1] use this to study ordinary cohomology, taking R to be the Eilenberg-MacLane spectrum Hk associated to a commutative ring k. Such Eilenberg-MacLane spectra are commutative S-algebras by multiplicative infinite loop space theory [8] and the theory of [3].

1. Formal results on change of universe

All change of universe functors are obtained as examples of twisted half-smash products. Let U and U′ be G-universes and let ℐ (U, U′) be the space of linear isometries U ⟶ U′, with G acting by conjugation. Let A be a G-space and let α : A ⟶ ℐ (U, U′) be a G-map. The twisted half-smash product is a functor G𝒮 U ⟶ G𝒮 U′, written A ⋉ E on objects; this is an abuse of notation since the functor depends on α and not just A. Different α give rise to equivalent functors on passage to derived categories. See [3, I§§2,3] for a summary of the properties of this functor and [7, Ch. VI] for details of its construction.

A G-linear isometry f : U ⟶ U′ may be regarded as a G-map {∗} ⟶ ℐ (U, U′). The corresponding twisted half-smash product functor is denoted f_{∗} : G𝒮 U ⟶ G𝒮 U′. The existence of such a G-linear isometry ensures that ℐ (U, U′) is a Gcontractible G-space [7, II.1.5], and it follows that different choices of f give rise to equivalent functors on passage to stable homotopy categories. These are the standard change of universe functors. However, there is a more canonical choice, namely ℐ (U, U′) ⋉ E. For reasonable G-spectra, namely tame ones, the G-equivalence {∗} ⟶ ℐ (U, U′) determined by f induces a homotopy equivalence of G-spectra f_{∗}E ⟶ ℐ (U, U′) ⋉ E [3, I.2.5].

We now sketch the equivariant versions of the basic definitions of [3, I, II]. The details are identical with those given there. Fix a universe U. There is a monad 𝕃 on the category G𝒮 U such that 𝕃E = ℐ (U, U) ⋉ E. The unit and product of 𝕃 are induced by the inclusion of the identity isometry in ℐ (U, U) and by the composition product ℐ (U, U) × ℐ (U, U) ⟶ ℐ (U, U). An 𝕃-spectrum is an algebra over the monad 𝕃, and a map of 𝕃-spectra is a map of 𝕃-algebras.

Given two 𝕃-spectra M and N, their operadic smash product is the 𝕃-spectrum

M ∧_{ℒ} N = ℐ (U ⊕ U, U) ⋉_{ℐ (U,U)×ℐ (U,U)} (M ∧ N).

Here ∧ on the right is the external smash product G𝒮 U × G𝒮 U ⟶ G𝒮 (U ⊕ U). The construction is made precise by a coequalizer diagram based on the evident right action of ℐ (U, U)×ℐ (U, U) on ℐ (U ⊕U, U) and a left action of ℐ (U, U)×ℐ (U, U) on M ∧ N induced by the actions of ℐ (U, U) on M and N; the details are just like those in the following definition. The term “operadic” refers to the operad ℒ in the category of G-spaces whose jth G-space is ℒ (j) = ℐ (U^{j}, U). The operadic smash product is associative and commutative, and there is a natural weak equivalence λ : SU ∧_{ℒ} M ⟶ M.

An SU-module is an 𝕃-spectrum M for which λ is an isomorphism, and a map of SU-modules is a map of 𝕃-spectra between them. For SU-modules M and N, M ∧_{ℒ} N is again an SU-module and is denoted M ∧_{SU} N. This smash product of SU-modules is commutative, associative, and unital, with unit SU. That is, the category Gℳ_{SU} of SU-modules is symmetric monoidal under its smash product.

We obtain the homotopy category hGℳ_{SU} by identifying homotopic maps of SUmodules, and we obtain the derived category G𝒟_{SU} by formally inverting those maps of SU-modules whose underlying maps of G-spectra are weak equivalences. This is made rigorous by CW-approximation. We emphasize that all of this applies to any G-universe U.

With motivation exactly as in the definition of the smash product over ℒ, this leads us inexorably to the following version of our change of universe functors.

Definition 1.1. Fix universes U and U′ and write 𝕃 and 𝕃′ for the respective monads in G𝒮 U and G𝒮 U′ and ℒ and ℒ′ for the respective operads of G-spaces. For an 𝕃-spectrum M, define an 𝕃′-spectrum I_{U}^{U′}M by

I_{U}^{U′}M = ℐ (U, U′) ⋉_{ℐ (U,U)} M.

That is, I_{U}^{U′}M is the coequalizer displayed in the diagram

[figure: ℐ (U, U′) ⋉ (ℐ (U, U) ⋉ M); γ⋉id; id ⋉ξ; ℐ (U, U′) ⋉ M; I_{U}^{U′}M.]

Here ξ : ℐ (U, U) ⋉ M ⟶ M is the given action of 𝕃 on M. We regard ℐ (U, U′) × ℐ (U, U) as a space over ℐ (U, U′) via the composition product

γ : ℐ (U, U′) × ℐ (U, U) ⟶ ℐ (U, U′),

and there results a natural isomorphism

ℐ (U, U′) ⋉ (ℐ (U, U) ⋉ M) ≅ (ℐ (U, U′) × ℐ (U, U)) ⋉ M.

This makes sense of the map γ ⋉ id in the diagram. The required left action of ℐ (U′, U′) on I_{U}^{U′}M is induced by the composition product

γ : ℐ (U′, U′) × ℐ (U, U′) ⟶ ℐ (U, U′),

which induces a natural map of coequalizer diagrams on passage to twisted half-smash products.

We need the following easy observation, in which we do not restrict to universes.

Lemma 1.2. Let U, U′, and U′′ be G-inner product spaces. Assume that either U is nonequivariantly isomorphic to U′ or U′ is nonequivariantly isomorphic to U′′. Then the diagram

[figure: ℐ (U′, U′′) × ℐ (U′, U′) × ℐ (U, U′); γ×id; id ×γ; ℐ (U′, U′′) × ℐ (U, U′); γ; ℐ (U, U′′)]

is a split coequalizer of spaces and therefore a coequalizer of G-spaces. Thus

ℐ (U, U′′) ≅ ℐ (U′, U′′) ×_{ℐ (U′,U′)} ℐ (U, U′).

Proof. Define maps

h : ℐ (U, U′′) ⟶ ℐ (U′, U′′) × ℐ (U, U′)

and

k : ℐ (U′, U′′) × ℐ (U, U′) ⟶ ℐ (U′, U′′) × ℐ (U′, U′) × ℐ (U, U′)

as follows. If s : U ⟶ U′ is an isomorphism, define

h(f) = (f ∘ s^{−1}, s) and k(g′, g) = (g′, g ∘ s^{−1}, s).

If t : U′ ⟶ U′′ is an isomorphism, define

h(f) = (t, t^{−1} ∘ f) and k(g′, g) = (t, t^{−1} ∘ g′, g).

In the first case, γ ∘ h = id, (id ×γ) ∘ k = id, and (γ × id) ∘ k = h ∘ γ. In the second case, γ ∘ h = id, (γ × id) ∘ k = id, and (id ×γ) ∘ k = h ∘ γ. In either case, this proves that we have a split coequalizer of spaces. Since coequalizers of G-spaces are created in the underlying category of spaces, it follows that we have a coequalizer of G-spaces, although not necessarily a split one.

Let G𝒮 U[𝕃] denote the category of 𝕃-spectra. Write Σ^{∞}U for the suspension G-spectrum functor from the category G𝒯 of based G-spaces to G𝒮 U. This functor takes values in G𝒮 U[𝕃] and in fact in Gℳ_{SU} [3, II.1.2].

Proposition 1.3. Let U, U′, and U′′ be G-universes. Consider the functors

I_{U}^{U′} : G𝒮 U[𝕃] ⟶ G𝒮 U′[𝕃′] and Σ^{∞}U : G𝒯 ⟶ G𝒮 U[𝕃].

(i) I_{U}^{U′} ∘ Σ^{∞}U is naturally isomorphic to Σ^{∞}U′.

(ii) I_{U′}^{U′′} ∘ I_{U}^{U′} is naturally isomorphic to I_{U}^{U′′}.

(iii) I_{U}^{U} is naturally isomorphic to the identity functor. Therefore the functor I_{U}^{U′} is an equivalence of categories with inverse I_{U′}^{U}. Moreover, the functor I_{U}^{U′} is continuous and satisfies I_{U}^{U′}(M ∧ X) ≅ (I_{U}^{U′}M) ∧ X for 𝕃-spectra M and based G-spaces X. In particular, it is homotopy preserving, and I_{U}^{U′} and I_{U′}^{U} induce inverse equivalences of homotopy categories.

Proof. For (i), we view G𝒯 as the category of “G-spectra indexed on {0}” and have that (Σ^{∞}U)(X) is isomorphic to ℐ ({0}, U) ⋉ X. By [3, I.2.2(ii)] and the previous lemma, we find

(I_{U}^{U′}∘Σ^{∞}U)(X) ≅ [ℐ (U, U′)×_{ℐ (U,U)} ℐ ({0}, U)]⋉X ≅ ℐ ({0}, U′)⋉X ≅ (Σ^{∞}U′)(X).

For (ii), [3, I.2.2(ii)] and the previous lemma give that

(I_{U′}^{U′′} ∘ I_{U}^{U′})(M) ≅ [ℐ (U′, U′′) ×_{ℐ (U′,U′)} ℐ (U, U′)] ⋉_{ℐ (U,U)} M ≅ I_{U}^{U′′}M.

Part (iii) is trivial since the relevant coequalizer splits. The topologies on the Hom sets of our categories are discussed in [3, VII§2]. The claimed continuity is easily checked, and commutation with smash products with spaces follows from the analogous property of twisted half-smash products [3, I.2.2(iv)].

These conclusions pass from 𝕃-spectra to Gℳ_{SU}-modules.

Proposition 1.4. The following statements hold.

(i) I_{U}^{U′}SU is canonically isomorphic to SU′.

(ii) For 𝕃-spectra M and N, there is a natural isomorphism

ω : I_{U}^{U′}(M ∧_{ℒ} N) ≅ (I_{U}^{U′}M) ∧_{ℒ′} (I_{U}^{U′}N).

(iii) The following diagram commutes for all 𝕃-spectra M:

[figure: I_{U}^{U′}(SU ∧_{ℒ} M); ω; I_{U}^{U′}λ; SU′ ∧_{ℒ′} (I_{U}^{U′}M); λ; I_{U}^{U′}M.]

(iv) M is an SU-module if and only if I_{U}^{U′}M is an SU′-module.

Therefore the functors I_{U}^{U′} and I_{U′}^{U} restrict to inverse equivalences of categories between Gℳ_{SU} and Gℳ_{SU′} that induce inverse equivalences of categories between hGℳ_{SU} and hGℳ_{SU′}.

Proof. The isomorphism in (i) is obtained by applying (i) of the previous proposition to the space S^{0}. For (ii), [3, I.2.2(ii)], Lemma 1.2, and a slight generalization of its analog [3, I.5.4] give that

I_{U}^{U′}(M ∧_{ℒ} N)
= ℐ (U, U′) ⋉_{ℐ (U,U)} [ℐ (U^{2}, U) ⋉_{ℐ (U,U)^{2}} (M ∧ N)]
≅ [ℐ (U, U′) ×_{ℐ (U,U)} ℐ (U^{2}, U)] ⋉_{ℐ (U,U)^{2}} (M ∧ N)
≅ ℐ (U^{2}, U′) ⋉_{ℐ (U,U)^{2}} (M ∧ N)

and

(I_{U}^{U′}M) ∧_{ℒ′} (I_{U}^{U′}N) = ℐ ((U′)^{2}, U′) ⋉_{ℐ (U′,U′)^{2}} (ℐ (U, U′) ⋉_{ℐ (U,U)} M) ∧ (ℐ (U, U′) ⋉_{ℐ (U,U)} N)

≅ [ℐ ((U′)^{2}, U′) ×_{ℐ (U′,U′)^{2}} ℐ (U, U′)^{2}] ⋉_{ℐ (U,U)^{2}} (M ∧ N)
≅ ℐ (U^{2}, U′) ⋉_{ℐ (U,U)^{2}} (M ∧ N).

Now (iii) is an exercise from the definition of λ in [3, I.8.3] and (iv) follows from (iii) since λ is an isomorphism if and only if I_{U}^{U′}λ is an isomorphism.

Since the smash product of SU-modules is their smash product as 𝕃-spectra, the previous result has the following immediate consequence. Theorem 0.1 is the special case that is obtained by replacing the pair (U, U′) by the pair (U^{G}, U).

Theorem 1.5. The functor I_{U}^{U′} : Gℳ_{SU} ⟶ Gℳ_{SU′} is monoidal. If R is an SUalgebra and M is an R-module, then I_{U}^{U′}R is an SU′-algebra and I_{U}^{U′}M is an I_{U}^{U′}Rmodule.

2. Homotopical results on change of universe

Obviously, the functor I_{U}^{U′} cannot preserve weak equivalences in general. To obtain a functor G𝒟_{SU} ⟶ G𝒟_{SU′}, we take an SU-module M, construct a weak equivalence ΓM ⟶ M, where ΓM is a CW SU-module, and then apply the functor I_{U}^{U′}. This loses the good formal properties that we have just discussed. Moreover, we see that passage to derived categories in this fashion cannot preserve composition of functors. Since the functor on derived categories induced by I_{U}^{U} is still the identity, we cannot have I_{U′}^{U} ∘ I_{U}^{U′} ≅ I_{U}^{U} : G𝒟_{SU} ⟶ G𝒟_{SU}, as the case U′ = U^{G} makes abundantly clear. The point is that, on the level of derived categories, (I_{U′}^{U} ∘ I_{U}^{U′})(X) means I_{U′}^{U}(ΓI_{U}^{U′}ΓX), and, because of the reapproximation in the middle, this need not and generally will not be equivalent to I_{U}^{U}ΓX.

Suppose given a G-linear isometry f : U ⟶ U′ and consider the composite

α : f_{∗}M ⟶ ℐ (U, U′) ⋉ M ⟶ ℐ (U, U′) ⋉_{ℐ (U,U)} M = I_{U}^{U′}M,

where M is an 𝕃-spectrum. As we have already pointed out, the first arrow is a homotopy equivalence when M is tame, for example, when M has the homotopy type of a G-CW spectrum. We shall prove that the second arrow, and therefore α, is a homotopy equivalence for a large class of SU-modules.

Fix U for the moment. For a G-spectrum X and j ≥ 0, define an 𝕃-spectrum D_{j}X by

D_{j}X = ℒ (j) ⋉_{Σ_{j}} X^{j}, where ℒ (j) = ℐ (U^{j}, U).

By convention, D_{0}X = SU for any X; D_{1}X = ℐ (U, U) ⋉ X is the free 𝕃-spectrum generated by X. If M is a CW 𝕃-spectrum, then M is homotopy equivalent to 𝕃X for some G-CW spectrum X [3, I.4.7]. The functor SU ∧_{ℒ} (?) converts 𝕃-spectra to SU-modules [3, II.1.3]. If M is a CW SU-module, then M is homotopy equivalent to SU ∧_{ℒ} (𝕃X) for some G-CW spectrum X [3, II.1.9].

We repeat the following definition from [3, VII.6.4].

Definition 2.1. Let ℰ_{SU} be the collection of SU-modules of the form

SU ∧_{ℒ} D_{j}X,

where X is any G-spectrum of the homotopy type of a G-CW spectrum and j ≥ 0. Let ℰ̄_{SU} be the closure of ℰ_{SU} under finite ∧_{SU}-products, wedges, pushouts along cofibrations, colimits of countable sequences of cofibrations, and homotopy equivalences.

Clearly all SU-modules of the homotopy types of G-CW SU-modules are in ℰ̄_{SU}. This class also contains an SU-algebra weakly equivalent to any given SU-algebra. In fact, there are Quillen model structures on the categories of SU-algebras and of commutative SU-algebras. Every SU-algebra or commutative SU-algebra is weakly equivalent to one which is q-cofibrant in the relevant model structure. The following result is part of the equivariant version of [3, VII.6.5].

Theorem 2.2. The underlying SU-module of a q-cofibrant SU-algebra or commutative SU-algebra is in ℰ̄_{SU}. Therefore any cell module over a q-cofibrant SU-algebra or q-cofibrant commutative SU-algebra is in ℰ̄_{SU}.

The following result is part of the equivariant version of [3, VII.6.6].

Proposition 2.3. The underlying G-spectrum of any SU-module in ℰ̄_{SU} has the homotopy type of a G-CW spectrum.

Returning to consideration of change of universe and again recording universes in our notations, we have the following theorem. It states that, for SU-modules in ℰ̄_{SU}, the change of universe functor f_{∗} and the functor I_{U}^{U′} are weakly equivalent and can therefore be used interchangeably on passage to derived categories. Theorem 0.2 is the special case that is obtained by replacing the pair (U, U′) by the pair (U^{G}, U) and taking f = i.

Theorem 2.4. Let f : U ⟶ U′ be a G-linear isometry. Then α : f_{∗}M ⟶ I_{U}^{U′}M is a homotopy equivalence of G-spectra indexed on U′ for every SU-module in ℰ̄_{SU}.

Proof. We begin by considering an 𝕃-spectrum D_{j}X for a G-CW spectrum X. In this case, we see by Lemma 1.2 that the second arrow in the definition of α is

ℐ (U, U′) ⋉ (ℐ (U^{j}, U) ⋉_{Σ_{j}} X^{j}) ≅ [ℐ (U, U′) × ℐ (U^{j}, U)] ⋉_{Σ_{j}} X^{j}
⟶ [ℐ (U, U′) ×_{ℐ (U,U)} ℐ (U^{j}, U)] ⋉_{Σ_{j}} X^{j} ≅ ℐ (U^{j}, U′) ⋉_{Σ_{j}} X^{j}.

The existence of the G-linear isometry f : U ⟶ U′ ensures that ℐ (U^{j}, U′) and ℐ (U, U′) × ℐ (U^{j}, U) are universal principal (G, Σ_{j})-bundles [7, II.2.11, VII.1.3] and thus that

γ : ℐ (U, U′) × ℐ (U^{j}, U) ⟶ ℐ (U^{j}, U′)

is a (G × Σ_{j})-homotopy equivalence. It follows from the equivariant version of [3, I.2.5] that the displayed map is a G-homotopy equivalence. We claim that we can replace D_{j}X by SU ∧_{ℒ} D_{j}X in this argument. In fact,

SU ∧_{ℒ} D_{j}X ≅ ℒ̃ (j) ⋉ X^{j},

where

ℒ̃ (j) = ℐ (U^{2}, U) ×_{ℐ (U,U)×ℐ (U,U)} ℐ (0, U) × ℐ (U^{j}, U).

By the equivariant version of [3, XI.2.2],

γ : ℒ̃ (j) ⟶ ℒ (j) = ℐ (U^{j}, U)

is a G × Σ_{j}-homotopy equivalence, and the claim follows. The theorem follows in view of the way that ℰ̄_{SU} is obtained from the SU ∧_{ℒ} D_{j}X.

References

1. D. J. Benson and J. P. C. Greenlees. Commutative algebra for cohomology rings of classifying spaces of compact Lie groups. Preprint. 1995.

2. A. D. Elmendorf, I. Kriz, M. A. Mandell, and J. P. May. Modern foundations for stable homotopy theory. In “A handbook of algebraic topology”, edited by I. M. James. Elsevier Science. 1995.

3. A. D. Elmendorf, I. Kriz, M. A. Mandell, and J. P. May. Rings, modules, and algebras in stable homotopy theory. Preprint 1995.

4. J. P. C. Greenlees and J. P. May. Generalized Tate cohomology. Memoirs Amer. Math. Soc. Vol 113, No 543. 1995.

5. J. P. C. Greenlees and J. P. May. Equivariant stable homotopy theory. In “A handbook of algebraic topology”, edited by I. M. James. Elsevier Science. 1995.

6. L. G. Lewis, Jr. Change of universe functors in equivariant stable homotopy theory. Preprint, 1995.

7. L. G. Lewis, Jr., J. P. May, and M. Steinberger (with contributions by J. E. McClure). Equivariant stable homotopy theory. Springer Lecture Notes in Mathematics Volume 1213. 1986.

8. J. P. May. Multiplicative infinite loop space theory. J. Pure and Applied Algebra 26(1983), 1–69.

Purdue University Calumet, Hammond IN 46323 E-mail address: aelmendo@math.purdue.edu

The University of Chicago, Chicago, IL 60637 E-mail address: may@math.uchicago.edu